Suppose based on a 36-observation time series sample (from January of year 1 to December of year 3) , the estimated model is y = 3.5 + 1.2t + 0.5t2, predict the y value for May of year 4.
Answer #1

time series sample (from January of year 1 to December of year 3)

y = 3.5 + 1.2t + 0.5t2

For May of year 4, we will use t = 36 +5 = 41

y = 3.5 + 1.2t + 0.5t2

y = 3.5 + 1.2*41 + 0.5*412

y = 3.5 + 49.2 + 840.5

y = 893.2